OpenStudy (anonymous):

What is 20% of 300 EXPLAIN

OpenStudy (aroub):

=20/100 *300 =60

OpenStudy (lalaly):

20%=20/100 = .2 300 * 0.2 = 60

OpenStudy (aroub):

llaannaa!!!

OpenStudy (lalaly):

aroubaaaaaaaaaaa;D

OpenStudy (anonymous):

@shanny try 20% of 300 in your head 1) move decimal on place left get 30 2) double 30 get 60

OpenStudy (anonymous):

Or 20/100 is a fifth..

OpenStudy (anonymous):

another example 20% of 70 1) 7 2)14

OpenStudy (anonymous):

another example 20% of 36 1) 3.6 2) 7.2

OpenStudy (anonymous):

thanks

OpenStudy (amistre64):

to explain: we want to divide up 300 pieces so that 100 people each get an equal share 300/100 = 3. If each person gets 3 pieces. If 20 of those people put all their pieces together into on pot; we would have 20% of 300, since each person holds 1% 3*20 = 60